# Resonance Reduction Against Adversarial Attacks

### v1

@@ Abstract
Resonance is a well-known phenomenon that happens in systems with second order dynamics. In this paper we address the fundamental question of making a network robust to signal being periodically pumped into it at or near a resonant frequency by an adversarial agent with the aim of saturating the network with the signal. Towards this goal, we develop the notion of network vulnerability, which is measured by the expected resonance amplitude on the network under a stochastically modeled adversarial attack. Assuming a second order dynamics model based on the network graph Laplacian matrix and a known stochastic model for the adversarial attack, we propose two methods for minimizing the network vulnerability that leverage the principle of eigenspectrum optimization. We provide extensive numerical results analyzing the effects of both methods.
@@
Network Graphs, Second Order Dynamics, Vulnerability Reduction
@@ I Introduction
While second order dynamics over networks has been studied in the past [1, 2, 3, 4], especially in context of power grids (since power transmission using alternating currents are described naturally using second-order dynamics), existing literature does not focus on controlling network parameters and topology for the purpose of filtering certain frequencies for prevention of resonance.
We consider networks whose dynamics are governed by second order differential equations where the coefficients are functions of the graph Laplacian matrix. Assuming an adversarial signal source that obeys a known stochastic model, we develop two methods (Network Graph Optimization and Auxiliary Graph Optimization) for optimizing the network structure to reduce signal resonance under the following conditions respectively: (i) network structure can be altered by modifying the edge weights (representing the connection strength between two network nodes), (ii) edge weights of the network cannot be modified directly, but an auxiliary network can be attached to the system.
The contributions of this paper are as follows:
  • We provide a second-order dynamics model of the network behavior under external forcing (source of the adversarial signal), that is consistent with the network topology (Section III).
@@
  • We propose two methods, namely Network Graph Optimization and Auxiliary Graph Optimization, both of which rely on the principle of eigenspectrum optimization (Sections IV and V).
@@ II Related Work
While first-order signal dynamics is most well-studied in context of networks [5, 7, 8], higher-order dynamics has also been studied. A second-order dynamics over a network is relevant, for example, in context of distributed power grids, electrical circuits and consensus in such networks [3, 9], where the dynamics of alternating electrical current and voltage are naturally second order. The motion dynamics of mobile agents (e.g., robots) is often governed by Newtonian dynamics, which gives rise to second-order dynamics over a network of such agents [10]. Second order dynamics can also be used to model transmission of information on social networks where the transmissibility of a signal depends both on its amount (how widespread it is) and its rate of change (how “viral” it is). The properties of second-order dynamics over networks has been well-studied in the literature (see [1, 2] for example), and model reduction in the context of such dynamics has been investigated [3, 4]. However, existing literature does not focus on active control of network parameters and topology for the purpose of filtering certain frequencies for prevention of resonance.
In this paper we consider a general second-order dynamics over a network with external forcing. We particularly focus on developing methods for mitigating resonance attacks inflicted by an adversarial agent pumping oscillatory signal in a periodic manner at one or more vertices while trying to match a natural frequency of the network. To our knowledge, there has been no prior work on active detection and control of resonance in a general graphical network with a focus on increasing robustness of the network to adversarial attacks.
@@ III Motivation & Background
where, $K$ and $\Gamma$ are called the stiffness and damping matrices respectively that are consistent with the network topology (i.e., their $(k,j)$-th element is nonzero only if there exists an edge between the $k$-th and $j$-th vertices in the graph). The network is subject to an adversarial forcing specified by the forcing vector $\mathbf{f}$ (with its $k$-th element, $f_{k}$, being the amplitude of adversarial signal forced on the $k$-th vertex) and forcing frequency $\nu$.
@@
We assume that the adversarial agent tries to match its forcing frequency, $\nu$, with one of the natural frequencies of the system (one of the eigenvalues of $\Omega$), but, is subject to uncertainties, either due to an inability to precisely select the forcing frequency, or because of an imprecise knowledge of the natural frequencies of the system. In particular, we assume that $\nu$ is a stochastic variable with a probability density function dependent upon the natural frequencies of the system. We measure the vulnerability of a network against the attacks by the expected value of the squared 2-norm of the steady-state response, denoted as $\mathbb{E}_{{\nu,\!\mathbf{f}}}\left(\|\mathbf{x}_{s}\|^{2}_{2}\right)$.
The main objective of this work is to develop an approach that relies on the principle of eigenspectrum optimization to reduce the vulnerability of the network against adversarial attacks with a known stochastic model. We approach this problem in two different ways:
@@ IV Network Graph Optimization
Given an initial configuration of the main network specified via the graph ${G}$, the Network Graph Optimization (NGO), refers to the procedure of changing the main network graph’s weights and/or topology in such a way that the vulnerability of the network is minimized against the adversarial agent’s forcing behavior (forcing vector and frequency) obeying the stochastic model that will be explained in Sections IV-B and IV-A.
In this section, we derive an expression for the expected value of the squared 2-norm of the steady-state response (i.e. the vulnerability of the network), and formulate the eigenspectrum optimization problem to minimize the vulnerability of the network.
IV-A Stochastic Modeling of the Forcing Vector
@@
The proof of this lemma is deferred to Appendix -A for better readability. Using the result of this lemma and the identity $\|M\|_{F}^{2}=\operatorname{Tr}(M^{H}M)=\sum_{i=1}^{n}|\lambda_{i}(M)|^{2}$ (where $M^{H}$ denotes the conjugate transpose and $\lambda_{i}(M)$ denotes the $i$-th eigenvalue of $M$), the expected value of $\|\mathbf{x}_{s}\|^{2}_{2}$ (equation (3)) with respect to the random variable $\mathbf{f}$ is thus computed as follows:
@@
$=\frac{1}{n}\sum\limits_{k}\frac{1}{|-\nu^{2}+i2\nu\gamma\,\omega_{k}^{2}+\omega_{k}^{2}|^{2}}$
@@
where $\omega_{k}$ are the eigenvalues of $\Omega$.
As seen from equation (4), if $\nu$ matches one of the eigenvalues of $\Omega$, the corresponding term of the summation blows up when the damping, $\gamma$, is small, leading to resonance.
IV-B Stochastic Modeling of the Forcing Frequency
We assume that the adversarial agent has uncertain knowledge of the network (or equivalently precise knowledge of the network, but uncertainty/error in choosing a forcing frequency). This uncertainty/error manifests itself when the adversarial agent tries to pick a forcing frequency that matches with one of the natural frequencies of the network. We model this uncertainty by considering $\nu$ to be a random variable with a probability density function, $\rho$, constructed as a uniformly weighted sum of multiple Cauchy distributions [13] each of which are centered at the natural frequencies, $\{\omega_{j}\}_{j=1,\dots,n}$, with a constant spread of $h$:
The Cauchy distribution, as opposed to other probability distributions, allows the integral representing the expected value of $\|\mathbf{x}_{s}\|^{2}_{2}$ to converge. Figure 2 illustrates an example where three individual Cauchy distributions are summed up with uniform weights to obtain a composite probability distribution $\rho(\nu)$.
IV-C Derivation of the Optimization Objective
The expected value of $\|\mathbf{x}_{s}\|^{2}_{2}$ with respect to the random variables $\mathbf{f}$ and $\nu$ can then be calculated using equations (4) and (5) as follows:
[…]
$=\frac{h}{\pi n^{2}}\sum\limits_{k,j}g(\omega_{k},\omega_{j})$
@@ IV-C Derivation of the Optimization Objective
$g(\omega_{k},\omega_{j})=\int\limits_{-\infty}^{\infty}\frac{d\nu}{((\omega_{k}^{2}-\nu^{2})^{2}+(2\gamma\nu\omega_{k}^{2})^{2})((\omega_{j}-\nu)^{2}+h^{2})}$ (7)
If $\gamma$ is non-zero, the poles of the integrand above lie away from the real line on the complex plane, and hence a closed-form expression for the integral $g(\omega_{k},\omega_{j})$ can be obtained using the Residue theorem [14] by performing a contour integration over the real line and a semi-circular arc of radius $R\to\infty$ on the upper half of the complex plane (Figure 3).
@@
Assuming $\gamma\ll h$, we compute the roots of the quartic polynomial in $\nu$ in the denominator of the integrand in (7) using a symbolic algebra toolbox, and then apply the Residue theorem to obtain
$g(\omega_{k},\omega_{j})=\frac{\pi}{2\gamma}\frac{h^{2}+\omega_{k}^{2}+\omega_{j}^{2}}{\omega_{k}^{4}\left(h^{4}+2h^{2}(\omega_{k}^{2}+\omega_{j}^{2})+(\omega_{k}^{2}-\omega_{j}^{2})^{2}\right)}$ (8)
From (6) this leads to the following expression for the expected value of the $2$-norm of the steady-state amplitude:
$\mathbb{E}_{\mathbf{f},\nu}\left(\|\mathbf{x}_{s}\|^{2}_{2}\right)=\frac{h}{2\gamma n^{2}}\sum\limits_{k,j}\frac{h^{2}+\omega_{k}^{2}+\omega_{j}^{2}}{\omega_{k}^{4}\left(h^{4}+2h^{2}(\omega_{k}^{2}+\omega_{j}^{2})+(\omega_{k}^{2}-\omega_{j}^{2})^{2}\right)}$ (9)
The objective is to minimize this expected value of the $2$-norm of the steady-state amplitude, so as to mitigate the effects of resonance attacks on the network. We note that $\omega_{k}$ and $\omega_{j}$ are the eigenvalues of $\Omega=\sqrt{L+\epsilon I}$, where the Laplacian matrix, $L=D-A$, depends on the weights on the edges of the graph. Thus $\mathbb{E}_{\mathbf{f},\nu}\left(\|\mathbf{x}_{s}\|^{2}_{2}\right)$, as described in (9), is a function of the edge weights of the graph. We thus define the objective function, $J(\mathbf{w})=\mathbb{E}_{\mathbf{f},\nu}\left(\|\mathbf{x}_{s}\|^{2}_{2}\right)$ to be a function of the edge weight vector, $\mathbf{w}\in\mathbb{R}^{m}$ (where $m$ is the number of edges in the graph), given by (9). It can be checked that $J$ is in general a non-convex function.
IV-D Eigenspectrum Optimization of the Main Network Graph
@@
Figure 4: Histograms of the Laplacian matrix eigenvalues for the initial network graph ${G}$ and optimized network graph ${G}^{*}$. The initial network is modeled by a complete graph, whose edge weights are perturbed away from a uniform distribution by a small amount. The corresponding eigenspectrum (on the left) is peaky, whereas as a result of the eigenspectrum optimization, the eigenvalue spectrum (on the right) has become flatter.
@@ V Auxiliary Graph Optimization
In this section, we first reformulate the dynamics equations and the definition of vulnerability based on the combined network (main network $+$ auxiliary network). Then, we derive the corresponding objective function and formulate the eigenspectrum optimization problem to minimize the vulnerability of the main network.
V-A Reformulation of Combined Dynamics

### v2

# Spectrum Optimization of Dynamic Networks for

@@ Abstract
Resonance is a well-known phenomenon that happens in systems with second order dynamics. In this paper we address the fundamental question of making a network robust to signal being periodically pumped into it at or near a resonant frequency by an adversarial agent with the aim of saturating the network with the signal. Towards this goal, we develop the notion of network vulnerability, which is measured by the expected resonance amplitude on the network under a stochastically modeled adversarial attack. Assuming a second order dynamics model based on the network graph Laplacian matrix and a known stochastic model for the adversarial attack, we propose two methods for minimizing the network vulnerability through optimization of the spectrum of the network graph. We provide extensive numerical results analyzing the effects of both methods.
@@
Second-order Signal Dynamics on Graphs, Optimization, Algebraic/Geometric Methods, Network Vulnerability Reduction
@@ I Introduction
While second order dynamics over networks has been studied in the past [1, 2, 3, 4], especially in context of power grids (since power transmission using alternating currents are described naturally using second-order dynamics), existing literature does not focus on controlling network parameters and topology for the purpose of mitigation of resonance.
We consider networks whose dynamics are governed by second order differential equations where the coefficients are functions of the graph Laplacian matrix. Assuming an adversarial signal source that obeys a known stochastic model, we develop two methods (Network Graph Optimization and Auxiliary Graph Optimization) for optimizing the network structure to reduce signal resonance under the following conditions respectively: (i) network structure can be altered by modifying the edge weights (representing the connection strength between two network nodes), (ii) edge weights of the network cannot be modified directly, but an auxiliary network can be attached to the system. The contributions of this paper are as follows:
  • We provide a second-order dynamics model for signal transmission over a network under external forcing (source of the adversarial signal), that is consistent with the network topology (Section III).
@@
  • We propose two methods, namely Network Graph Optimization and Auxiliary Graph Optimization, both of which rely on the principle of graph spectrum optimization (Sections IV and V).
@@ II Related Work
While first-order signal dynamics is most well-studied in context of networks [5, 7, 8], higher-order dynamics has also been studied. A second-order dynamics over a network is relevant, for example, in context of distributed power grids, electrical circuits and consensus in such networks [3, 9, 10], where the dynamics of alternating electrical current and voltage are naturally second order. The motion dynamics of mobile agents (e.g., robots) is often governed by Newtonian dynamics, which gives rise to second-order dynamics over a network of such agents [11]. Second order dynamics can also be used to model transmission of information on social networks where the transmissibility of a signal depends both on its amount (how widespread it is) and its rate of change (how “viral” it is). The properties of second-order dynamics over networks have been well-studied in the literature (see [1, 2] for example), and model reduction in the context of such dynamics has been investigated [3, 4]. However, existing literature does not focus on active control of network parameters and topology for the purpose of prevention of resonance.
Optimization of the spectrum of the Laplacian matrix in order to affect the connectivity of a network has also been extensively studied [12, 13, 14]. However, most often, such optimization problems focus on the network connectivity in general, without explicitly addressing performance of a second-order signal dynamics over the network.
In this paper we consider a general second-order dynamics over a network with external forcing. We particularly focus on developing methods for mitigating resonance attacks inflicted by an adversarial agent pumping oscillatory signal in a periodic manner at one or more vertices while trying to match a natural frequency of the network. To our knowledge, there has been no prior work on control of resonance in a general graphical network with a focus on increasing robustness of the network to adversarial attacks.
@@ III Motivation & Background
where, $K$ and $\Gamma$ are called the stiffness and damping matrices respectively that are consistent with the network topology (i.e., their $(k,j)$-th element is nonzero only if there exists an edge between the $k$-th and $j$-th vertices in the graph). The network is subject to an adversarial forcing vector $\mathbf{f}$ (with its $k$-th element, $f_{k}$, being the amplitude of adversarial signal forced on the $k$-th vertex) and forcing frequency $\nu$.
@@
We assume that the adversarial agent tries to match its forcing frequency, $\nu$, with one of the natural frequencies of the system (one of the eigenvalues of $\Omega$), but, is subject to uncertainties, either due to an inability to precisely select the forcing frequency, or because of an imprecise knowledge of the natural frequencies of the system. In particular, we assume that $\nu$ is a stochastic variable with a probability density function dependent upon the natural frequencies of the system.
Definition 1 (Network Vulnerability to Adversarial Resonance Attack).
We define the network vulnerability to adversarial resonance attack to be the expected value of the squared 2-norm of the steady-state response, denoted as $\mathbb{E}_{{\nu,\!\mathbf{f}}}\left(\|\mathbf{x}_{s}\|^{2}_{2}\right)$
The main objective of this work is to develop approaches for optimization of the spectrum of the network graph (i.e., the spectrum of the Laplacian matrix, or equivalently, the spectrum of $\Omega^{2}$) to reduce the vulnerability of the network against adversarial resonance attacks with a known stochastic model. We approach this problem in two different ways:
@@ IV Network Graph Optimization
Given an initial configuration of the main network specified via the graph ${G}$, the Network Graph Optimization, refers to the procedure of optimizing the main network graph’s weights and/or topology in such a way that the vulnerability of the network is minimized against the adversarial agent’s forcing behavior (forcing vector and frequency) obeying the stochastic model that will be explained in Section IV-A.
In this section, we formulate the spectrum optimization problem to minimize the vulnerability of the network (i.e., the expected value of the squared 2-norm of the steady state response).
IV-A Stochastic Model of the Adversarial Forcing
@@
We assume that the adversarial agent has uncertain knowledge of the network (or equivalently precise knowledge of the network, but uncertainty/error in choosing a forcing frequency). This uncertainty/error manifests itself when the adversarial agent tries to pick a forcing frequency that matches one of the natural frequencies of the network. We model this uncertainty by considering $\nu$ to be a random variable whose probability density function, $\rho$, is a uniformly weighted sum of multiple Cauchy distributions [17], each of which are centered at the natural frequencies, $\{\omega_{j}\}_{j=1,\dots,n}$, with a constant spread of $h$:
The Cauchy distribution, as opposed to other probability distributions, allows the integral representing the expected value of $\|\mathbf{x}_{s}\|^{2}_{2}$ to be efficiently computed. Figure 2 illustrates an example where three individual Cauchy distributions are summed up with uniform weights to obtain a composite probability distribution $\rho(\nu)$.
IV-B Network Vulnerability
Following proposition computes the network vulnerability in terms of the spectrum of the network.
Proposition 1 (Network vulnerability).
If $\gamma<<h$, then the network vulnerability (i.e., the expected value of the 2-norm of the steady state amplitude) is given by:
$\mathbb{E}_{\mathbf{f},\nu}\left(\|\mathbf{x}_{s}\|^{2}_{2}\right)=\frac{h}{2\gamma n^{2}}\sum\limits_{k,j}\frac{h^{2}+\omega_{k}^{2}+\omega_{j}^{2}}{\omega_{k}^{4}\left(h^{4}+2h^{2}(\omega_{k}^{2}+\omega_{j}^{2})+(\omega_{k}^{2}-\omega_{j}^{2})^{2}\right)}$
where $\omega_{k}$ and $\omega_{j}$ are the eigenvalues of $\Omega$.
In order to prove this result we need the following lemmas:
@@ IV-B Network Vulnerability
The proof of the above lemma is deferred to Appendix -A for better readability.
Lemma 2.
If $M_{1}$ and $M_{2}$ are real symmetric matrices that commute, then $\|(M_{1}+iM_{2})^{-1}\|_{F}^{2}=\sum_{j=1}^{n}\frac{1}{\lambda_{j}(M_{1})^{2}+\lambda_{j}(M_{2})^{2}}$, where $\lambda_{j}(M)$ denotes the $j$-th eigenvalue of $M$.
The above lemma follows from the definition of the Frobenius norm, $\|M\|_{F}=\sqrt{\operatorname{Tr}(M^{*}M)}$ (where $M^{*}$ denotes the conjugate transpose of $M$).
Proof.
of Proposition 1: The expected value of $\|\mathbf{x}_{s}\|^{2}_{2}$ with respect to the random variables $\mathbf{f}$ and $\nu$ is calculated as follows:
[…]
From Lemma 1 and 2, we have:
@@
Substituting equation (6) into equation (5),
$\mathbb{E}_{\mathbf{f},\nu}\left(\|\mathbf{x}_{s}\|^{2}_{2}\right)=\frac{h}{\pi n^{2}}\sum\limits_{k,j}g(\omega_{k}^{2},\omega_{j}^{2})$ (7)
@@
$g(\omega_{k}^{2},\omega_{j}^{2})=\int\limits_{-\infty}^{\infty}\frac{d\nu}{((\omega_{k}^{2}-\nu^{2})^{2}+(2\gamma\nu\omega_{k}^{2})^{2})((\omega_{j}-\nu)^{2}+h^{2})}$ (8)
Since $\gamma$ is non-zero, the poles of the integrand above lie away from the real line on the complex plane, and hence a closed-form expression for the integral $g(\omega_{k}^{2},\omega_{j}^{2})$ can be obtained using the Residue theorem [18] by performing a contour integration over the real line and a semi-circular arc of radius $R\to\infty$ on the upper half of the complex plane (Figure 3).
@@
Assuming $\gamma\ll h$, we can compute the roots of the quartic polynomial in $\nu$ in the denominator of the integrand in (8) using a symbolic algebra toolbox, and then apply the Residue theorem to obtain
$g(\omega_{k}^{2},\omega_{j}^{2})=\frac{\pi}{2\gamma}\frac{h^{2}+\omega_{k}^{2}+\omega_{j}^{2}}{\omega_{k}^{4}\left(h^{4}+2h^{2}(\omega_{k}^{2}+\omega_{j}^{2})+(\omega_{k}^{2}-\omega_{j}^{2})^{2}\right)}$ (9)
This proves the proposition. ∎
The objective is to minimize this expected value of the $2$-norm of the steady-state amplitude, so as to mitigate the effects of resonance attacks on the network. We note that $\omega_{k}$ and $\omega_{j}$ are the eigenvalues of $\Omega=\sqrt{L+\epsilon I}$, where the Laplacian matrix, $L=D-A$, depends on the weights on the edges of the graph. Thus $\mathbb{E}_{\mathbf{f},\nu}\left(\|\mathbf{x}_{s}\|^{2}_{2}\right)$, as described in Proposition 1, is a function of the edge weights of the graph. We thus define the objective function, $J(\mathbf{w})=\mathbb{E}_{\mathbf{f},\nu}\left(\|\mathbf{x}_{s}\|^{2}_{2}\right)$ to be a function of the edge weight vector, $\mathbf{w}\in\mathbb{R}^{m}$ (where $m$ is the number of edges in the graph). It can be checked that $J$ is in general a non-convex function. However, if $h\rightarrow 0$, it can be indeed shown that $J$ is convex in the edge weights.
Proposition 2.
For a sufficiently large value of $h$, $J(\mathbf{w})$ is convex.
Proof.
Sketch: Define the symmetrized function $\widetilde{g}(\omega_{k}^{2},\omega_{j}^{2})=\frac{1}{2}\left(g(\omega_{k}^{2},\omega_{j}^{2})+g(\omega_{j}^{2},\omega_{k}^{2})\right)$ so that $J(\mathbf{w})=\frac{h}{\pi n^{2}}\sum_{k,j}\widetilde{g}(\omega_{k}^{2},\omega_{j}^{2})$.
Since $\{\omega_{j}^{2}\}_{j=1,2,\cdots,n}$ are eigenvalues of $\Omega^{2}=L+\epsilon I$, we can write $J(\mathbf{w})=\frac{h}{\pi n^{2}}\operatorname{Tr}(\widetilde{g}(L+\epsilon I,L+\epsilon I))$ (where $\widetilde{g}(M,N)$ refers to the matrix extension of the scalar function, $\widetilde{g}$ [19, 20]). It is known that the trace of the matrix extension of a scalar function inherits the convexity properties of the scalar function (see our technical report [19] for a detailed proof for the case of multi-variable scalar functions), and as a consequence of that, it’s sufficient to show that the function $\widetilde{g}$ is convex.
When $h$ is sufficiently large (compared to the eigenvalues of $L$), the function $\widetilde{g}$ becomes $\widetilde{g}(x,y)=\frac{\pi}{4\gamma}\frac{h^{2}+x+y}{h^{4}+2h^{2}(x+y)+(x-y)^{2}}\left(\frac{1}{x^{2}}+\frac{1}{y^{2}}\right)\simeq\frac{\pi}{4\gamma h^{2}}\left(\frac{1}{x^{2}}+\frac{1}{y^{2}}\right)$. It s easy to show that this function in convex in $\mathbb{R}_{+}^{2}$ (a direct computation of the Hessian shows that its eigenvalues are positive). This proves the proposition. ∎
As a consequence of the above proposition, while $J(\mathbf{w})$ may not be strictly convex for all values of $h$, when $h$ is large (corresponding to high uncertainty in the adversarial agent’s ability to choose/apply a forcing that matches a natural frequency of the graph), the objective is indeed concvex.
IV-C Spectrum Optimization of the Main Network Graph
@@
Figure 4: Histograms of the Laplacian matrix eigenvalues for the initial network graph ${G}$ and optimized network graph ${G}^{*}$. The initial network is modeled by a complete graph, whose edge weights are perturbed away from a uniform distribution by a small amount. The corresponding spectrum (on the left) is peaky, whereas as a result of the spectrum optimization, the spectrum (on the right) has become flatter.
@@ V Auxiliary Graph Optimization
In this section, we first reformulate the dynamics equations and the definition of vulnerability based on the combined network (main network $+$ auxiliary network). Then, we derive the corresponding objective function and formulate the spectrum optimization problem to minimize the vulnerability of the main network.
V-A Formulation of Combined Dynamics
@@
V-B Network Vulnerability with Attached Auxiliary Network
Following proposition gives the vulnerability of a network to which we attach the auxiliary network.
Proposition 3 (Network vulnerability with attached auxiliary network).
If ${\Omega}$ and ${\widetilde{\Omega}}$ are simultaneously diagonalizable, then the vulnerability of a network to which an auxiliary network is attached is given by:
[…]
where $\rho_{\omega_{j}}(\nu)=\frac{h/\pi}{({\omega}_{j}-\nu)^{2}+h^{2}}$, and,
$s_{k}(\nu)=\frac{1}{-\nu^{2}+i\nu 2{\gamma}{\omega}_{k}^{2}+{\omega}_{k}^{2}+{c}-\frac{{c}^{2}}{-\nu^{2}+i\nu 2{\widetilde{\gamma}}{\widetilde{\omega}}_{k}^{2}+{\widetilde{\omega}}_{k}^{2}+{c}}}$
with ${\widetilde{\omega}}_{k}$ denoting the $k$-th eigenvalue of ${\widetilde{\Omega}}$ and $\overline{s}_{k}$ denoting the complex conjugate of $s_{k}$.
Proof.